# A Foundational Individual Mobility Prediction Model based on Open-Source Large Language Models

### v2

@@ 3.4 Instruction Data Generation
Since most current LLMs are initially designed for dealing with various natural language tasks, there is a non-trivial mismatch between the original training objective of LLMs and our studied individual mobility prediction problem. To address the mismatch, we adopt the instruction tuning approach, which is commonly applied to enhance the capabilities and controllability of LLMs in other domain-specific tasks (Zhang et al., 2023c). The key to the approach is to construct an instruction-tuning dataset for LLMs’ fine-tuning. Each sample in the dataset can be concisely represented as an $(Instruction,Output)$ pair. $Instruction$ is a kind of prompt for instructing LLMs to specialize in the studied problem during the fine-tuning process, as shown in Figure 3. The $Output$ denotes the desired output that follows the $Instruction$.
In practical application scenarios, users have different styles in writing prompts. The LLMs fine-tuned by single-style instructions have difficulty adapting to varying prompts, which probably leads to model performance reduction. To address this issue, we adopt an Alpaca-style prompt (Taori et al., 2023) to distill different expressions describing individual mobility prediction tasks to generate sufficient multi-style task instructions as the examples shown in Figure 2. Typically, distillation imparts knowledge and cognitive abilities from a competent teacher model to a simpler and more computationally efficient student model (Zhang et al., 2023c).
@@
Figure 4 shows the CoT-based prompt for generating multi-style instructions with three main parts (see Appendix for generated examples). First, the LLMs should know how to generate diverse instructions related to time series prediction. Then, the requirements assist LLMs to work like professional mobility researchers to focus on the goal of human mobility prediction. In particular, it is necessary to limit the generation scope to text and obligatorily consider the generated instruction structure in the last three requirements. Such structural information can guide LLMs to generate CoT-based task instructions, which have been demonstrated to improve LLMs’ mobility prediction performance by previous studies (Wang et al., 2023b; Liang et al., 2024; Feng et al., 2024b; Qin et al., 2025).
To ensure the data authenticity, only the task definition, data description, and thinking guidance in Figure 3 are generated with diverse styles to get semi-complete instructions in Figure 2. Then it combines with a piece of real-world mobility data from the database to get an $Instruction$. To align with the required prediction results of the task prompt, the corresponding $Output$ given an $Instruction$ is set as $\{prediction:place\_id\}$ where the $place\_id$ is the ground truth of the next location or origin / destination station defined in the Problem Definition.
@@
Figure 3: The example of instruction data. For the data input, $<$history$>$ and $<$context$>$ are the user’s mobility sequences respectively, containing long-term and recent stays or trips. $<$target$>$ includes the information about current states except for location, e.g., the tap-out station and time of the last trip.
@@ 3.5 Parameter Efficient Fine-Tuning
Fine-tuning is a widely adopted approach for adapting LLMs to domain-specific datasets. It leverages the pre-trained knowledge embedded in LLMs to enhance task-specific performance while requiring relatively limited data and computational resources (Parthasarathy et al., 2024). However, conventional approaches such as full-parameter fine-tuning demand substantial Graphics Processing Unit (GPU) memory and computational overhead, making them impractical for use on resource-constrained devices or small-scale workstations. In contrast, parameter-efficient fine-tuning (PEFT) methods significantly reduce memory requirements while maintaining high adaptation performance (Han et al., 2024). One of the most prominent PEFT techniques is Low-Rank Adaptation (LoRA) (Hu et al., 2021), which has been extensively used to fine-tune LLMs in mobility-related applications (Li et al., 2024; Gong et al., 2024).
@@
where $h$ denotes the output embedding aligned with mobility-related semantics in $X$. Here, $\alpha$ is a tunable scaling factor used to control the contribution of the low-rank update.
Instead of updating the full weight matrix, LoRA introduces a low-rank correction term $W_{1}W_{2}$ on top of $W_{0}$, allowing the model to efficiently adapt without retraining all parameters. Since the number of parameters in $W_{1}$ and $W_{2}$ is orders of magnitude smaller than in $W_{0}$, this approach significantly reduces GPU memory requirements. Empirical studies report that LoRA can achieve up to a 3 times reduction in memory usage while maintaining competitive performance (Hu et al., 2021), making it well-suited for fine-tuning LLMs on mobility datasets under limited resources.
@@ 4.1 Data Description and Performance metrics
Dataset Type User Number Test Samples District Task Type
Geolife GPS trajectory 45 3459 Beijing task 1
FSQ-NYC check-in 535 12141 New York task 2
FSQ-TYK check-in 450 11250 Tokyo task 2
@@
task 2
HK-ORI AFC 500 5260 Hong Kong task 3
HK-DEST AFC 500 5260 Hong Kong task 4
HZ-ORI AFC 1200 15296 Hangzhou task 3
HZ-DEST AFC 1200 15296 Hangzhou task 4
@@
    Weighted F1 measures the model’s performance in predicting frequently visited locations to avoid the issue of imbalance of individual’s visited locations, which is denoted as the abbreviation $F1$ in the following parts.
    $F1=\frac{2TP}{2TP+FP+FN},$ (4)
    where $TP$, $FP$ and $FN$ are respectively the number of true positive, false positive and false negative samples.
@@ 4.2 PEFT methods and parameters
There are two important hyperparameters in Equation (1) that have a significant impact on fine-tuning. They are the rank $r$ and scaling factor $\alpha$, which control the updated parameter size and the ratio of specific knowledge influencing the fine-tuning process. Their specific values used in the PEFT methods in Table 2 are referred to in the corresponding literature, considering the trade-off between GPU memory, fine-tuning time, and model performance. All PEFT methods used in this research are based on a pre-trained LLaMA-3.1-8B-Instruct  [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct except the LoftQ starting with a 4-bit and 64-rank checkpoint of LLaMA-3-8B-Instruct  [^5] https://huggingface.co/LoftQ/Meta-Llama-3-8B-Instruct-4bit-64rank provided. About 100K training samples are used for fine-tuning and the batch size is set to 3 for all the methods. All the model training and experimental evaluations are conducted on a workstation with Intel(R) Xeon(R) w7-3465X CPU, 4 NVIDIA RTX 6000 Ada GPUs, and 512-Gb memory.
@@ 4.3 Baselines
    DeepMove (Feng et al., 2018) uses recurrent networks with an attention mechanism to capture multi-scale temporal dependencies of mobility patterns.
@@ 4.5 Transferability Performance
The results are presented in Table 4, where the open-source MoBLLM model significantly outperforms the general-purpose commercial LLM. It demonstrates that an open-source LLM, i.e., LLaMA-3.1-8B-Instruct used in the case studies, can become an expert in different mobility prediction tasks by the proposed fine-tuning framework. Moreover, the MoBLLM shows a significant superiority in cost efficiency, mainly produced by generating instructions for training, which is about 60 times cheaper than a commercial LLM.
@@
4.6 Adaptation under Different Prediction Situations
To investigate the adaptation under different situations (events, incidents), we further estimate the model prediction performance in the Hong Kong AFC data (mainly due to the wide data availability for the test). As shown in Table 5, the adaptation datasets involve individual trips occurring in metro network change (MNC), policy intervention (PI), special events (SE), and metro incidents (MI). Besides, each of the four datasets can be divided into 2 subsets according to origin prediction (ORI) and destination prediction (DEST). All the results of validating adaptability are shown in Table 6, where the results of HK-ORI and HK-DEST in regular scenarios are also listed for comparison.
Overall, the MoBLLM model has robust performance in different prediction situations and significantly outperforms all the baseline models. The deep learning based models have huge performance reduction in PI-ORI and PI-DEST. Especially in PI-DEST, their performance reduces by over 10% compared with the performance in HK-DEST, while the LLM-based models only have a margin reduction of about 2%. The probable reason is that the PI datasets are much different from the HK-ORI and HK-DEST used for training the deep learning models. Deep learning models have low transferability, which leads to significant performance variation in situations where the input data is inconsistent with the training data to some extent.
We further analyze the robustness of the MoBLLM model in each prediction situation. In datasets MNC and PI that have fewer stations than the training data, the MoBLLM model performance has an average decrease by about 7.5% on ACC and F1 for origin prediction compared to the performance in HK-ORI. The MoBLLM model performs worse in PI than MNC for both origin and destination prediction because there exist policy interventions of price discounts that probably cause a part of passengers to change their travel behaviors. In the datasets SE and MI that have special events and incidents influencing the demands, the MoBLLM model has a margin performance reduction of about 3% for the origin prediction. It is because a small number of passengers may change their travel choices. For destination prediction on these two datasets, the MoBLLM even has better performance than the one on HK-DEST, which implies that the special events may not significantly influence the destination of existing trips. Therefore, the MoBLLM model has a good generalization ability to adapt to different prediction situations covering metro network change, policy disturbances, special events, and incidents.
Table 5: The descriptions of adaptation datasets
@@ 4.6 Adaptation under Different Prediction Situations
Table 6: Model performance comparison under different scenarios.
@@ 5 Conclusions
In the cases studies, six real-world mobility datasets from multiple cities are used to validate the performance of MoBLLM model. We systematically evaluate LoRA and its several advanced variants in fine-tuning to obtain the optimal MoBLLM model. Compared with the state-of-art baseline models on accuracy, the MoBLLM model averagely outperforms the best DL model (MHSA) by 50.44% and the LLM-Mob model based on commercial LLMs by 37.30%. Besides, the MoBLLM model also shows significant superiority in transferability and finance compared to the method based on commercial LLMs. In the adaptability evaluation on 4 trip datasets representing different conditional scenarios, the MoBLLM shows the robust performance and outperforms the best benchmark (MHSA) by 14.67% on average. These results of case studies comprehensively demonstrate the capacity of the MoBLLM model as the foundational individual mobility prediction model.
The MoBLLM model contributes to precise and robust provision of personalized information for transportation management, e.g., crowding and disruption management. However, note that the incident logs or event information are important for prediction. Future work will design the prompt (incorporating situational context) and explore the fine-tuning methods to further improve the MoBLLM’s adaptability for abnormal situations. Also, future validation of the MoBLLM model on data from more cities is valuable.

### v3

@@ 3.4 Instruction Data Generation
Since most current LLMs are initially designed for dealing with various natural language tasks, there is a non-trivial mismatch between the original training objective of LLMs and our studied individual mobility prediction problem. To address the mismatch, we adopt the instruction based fine-tuning approach, which is commonly applied to enhance the capabilities and controllability of LLMs in other domain-specific tasks (Zhang et al., 2023c). The key to the approach is to construct an instruction dataset for LLMs’ fine-tuning. Each sample in the dataset can be concisely represented as an $(Instruction,Output)$ pair. $Instruction$ is a kind of prompt for instructing LLMs to specialize in the studied problem during the fine-tuning process, as shown in Figure 3. The $Output$ denotes the desired output that follows the $Instruction$.
@@
In practical application scenarios, users have different styles in writing prompts. The LLMs fine-tuned by single-style instructions have difficulty adapting to varying prompts, which probably leads to model performance reduction. To address this issue, we adopt an Alpaca-style prompt (Taori et al., 2023) to distill different expressions describing individual mobility prediction tasks to generate sufficient multi-style task instructions as the examples shown in Figure 2. Typically, distillation imparts knowledge and cognitive abilities from a competent teacher model to a simpler and more computationally efficient student model (Zhang et al., 2023c). Figure 4 shows the CoT-based prompt for generating multi-style instructions with three main parts (see A for generated examples). First, the LLMs should know how to generate diverse instructions related to time series prediction. Then, the requirements assist LLMs to work like professional mobility researchers to focus on the goal of human mobility prediction. In particular, it is necessary to limit the generation scope to text and obligatorily consider the generated instruction structure in the last three requirements. Such structural information can guide LLMs to generate CoT-based task instructions, which have been demonstrated to improve LLMs’ mobility prediction performance by previous studies (Wang et al., 2023b; Liang et al., 2024; Feng et al., 2024b; Qin et al., 2025). To ensure the data authenticity, only the task definition, data description, and thinking guidance in Figure 3 are generated with diverse styles to get semi-complete instructions in Figure 2. Then it combines with a piece of real-world mobility data from the database to get an $Instruction$. To align with the required prediction results of the task prompt, the corresponding $Output$ given an $Instruction$ is set as $\{prediction:place\_id\}$ where the place_idplace\_id is the ground truth of the next location or origin / destination station defined in the Problem Definition.
@@
Figure 3: The example of instruction data. For the data input, <<history>> and <<context>> are the user’s mobility sequences respectively, containing long-term and recent stays or trips. <<target>> includes the information about current states except for location, e.g., the tap-out station and time of the last trip.
@@ 3.5 Parameter Efficient Fine-Tuning
Fine-tuning is a widely adopted approach for adapting LLMs to domain-specific datasets. It leverages the pre-trained knowledge embedded in LLMs to enhance task-specific performance while requiring relatively limited data and computational resources (Parthasarathy et al., 2024). However, conventional approaches such as full-parameter fine-tuning demand substantial Graphics Processing Unit (GPU) memory and computational overhead, making them impractical for use on resource-constrained devices or small-scale workstations. In contrast, parameter-efficient fine-tuning (PEFT) methods significantly reduce memory requirements while maintaining high adaptation performance (Han et al., 2024). One of the most prominent PEFT techniques is Low-Rank Adaptation (LoRA) (Hu et al., 2022), which has been extensively used to fine-tune LLMs in mobility-related applications (Li et al., 2024; Gong et al., 2024).
@@
where $h$ denotes the output embedding aligned with mobility-related semantics in $X$. Here, $\alpha$ is a tunable scaling factor used to control the contribution of the low-rank update. Instead of updating the full weight matrix, LoRA introduces a low-rank correction term $W_{1}W_{2}$ on top of $W_{0}$, allowing the model to efficiently adapt without retraining all parameters. Since the number of parameters in $W_{1}$ and $W_{2}$ is orders of magnitude smaller than in $W_{0}$, this approach significantly reduces GPU memory requirements. Empirical studies report that LoRA can achieve up to a 3 times reduction in memory usage while maintaining competitive performance (Hu et al., 2022), making it well-suited for fine-tuning LLMs on mobility datasets under limited resources.
@@ 4.1 Data Description and Performance metrics
The MoBLLM model is obtained by following the pipeline in Figure 2 that fine-tunes a pre-trained LLaMA-3.1-8B-Instruct  [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct on the instruction dataset. We set MoBLLM’s temperature parameter to zero during inference to guarantee deterministic output across all case studies. In our transferability experiments, the training data were sampled from four datasets (Geolife, FSQ-NYC, HK-ORI and HK-DEST), while testing was conducted on four different datasets (FSQ-TYK, FSQ-Global, HZ-ORI, and HZ-DEST). As shown in Table 2, these datasets differ significantly in terms of data type (GPS trajectory, check-in, AFC), geographical context (Beijing, New York, Tokyo, Hong Kong, Hangzhou, and 415 global cities), and spatial coverage (ranging from dozens to over 100,000 unique locations). These differences suggest that the transfer experiments indeed involve substantial cross-dataset heterogeneity rather than trivial similarities.
@@
Dataset Type Users Test Samples District Locations Time Span Task Type
Geolife GPS trajectory 45 3459 Beijing 1145 2007/04-2012/08 task 1
FSQ-NYC Check-in 535 12141 New York 3752 2011/10-2012/02 task 2
FSQ-TYK Check-in 450 11250 Tokyo 15785 2012/04-2013/02 task 2
@@
112032 2012/04-2013/09 task 2
HK-ORI AFC 500 5260 Hong Kong 92 2018/01-2018/03 task 3
HK-DEST AFC 500 5260 Hong Kong 92 2018/01-2018/03 task 4
HZ-ORI AFC 1101 15296 Hangzhou 81 2019/01-2019/02 task 3
HZ-DEST AFC 1101 15296 Hangzhou 81 2019/01-2019/02 task 4
@@
    Weighted F1 measures the model’s performance in predicting frequently visited locations to avoid the issue of imbalance of individual’s visited locations. which is denoted as the abbreviation $F1$ in the following parts.
    $F1=\sum_{i\in D}\frac{s_{i}}{S}\cdot\frac{2TP_{i}}{2TP_{i}+FP_{i}+FN_{i}},$ (4)
    where $TP_{i}$, $FP_{i}$ and $FN_{i}$ are respectively the number of true positive, false positive and false negative samples given a ground truth location $l_{i}$. The $s_{i}$ represents samples of ground truth $l_{i}$ and the $S$ denotes the size of the dataset $D$. We use the F1 score weighted by the number of visits to emphasize the overall performance of the model in the most important locations.
@@ 4.2 PEFT methods and parameters
There are two important hyperparameters in Equation (1) that have a significant impact on fine-tuning. They are the rank $r$ and scaling factor $\alpha$, which control the updated parameter size and the ratio of specific knowledge influencing the fine-tuning process. Their specific values used in the PEFT methods in Table 3 are referred to in the corresponding literature, considering the trade-off between GPU memory, fine-tuning time, and model performance. All PEFT methods used in this research are based on a pre-trained LLaMA-3.1-8B-Instruct except the LoftQ starting with a 4-bit and 64-rank checkpoint of LLaMA-3-8B-Instruct  [^5] https://huggingface.co/LoftQ/Meta-Llama-3-8B-Instruct-4bit-64rank provided. About 100K training samples are used for fine-tuning and the batch size is set to 3 for all the PEFT methods following common practices in LoRA-based fine-tuning studies that adopt very small batch sizes due to GPU memory constraints (Li et al., 2024). All the model training and experimental evaluations are conducted on a workstation with Intel(R) Xeon(R) w7-3465X CPU, 4 NVIDIA RTX 6000 Ada GPUs, and 512-Gb memory.
@@ 4.3 Baselines
    DeepMove (1) uses recurrent networks with an attention mechanism to capture multi-scale temporal dependencies of mobility patterns.
@@ 4.5 Transferability Performance
The results are presented in Table 5, where the open-source MoBLLM model significantly outperforms the general-purpose commercial LLM. Both models perform worse on the check-in datasets than the AFC datasets since the former exhibit substantially higher sparsity in the set of possibly visited locations as shown in Table 2. This issue is particularly pronounced in FSQ-Global, which contains over 112,000 unique locations across 415 cities, resulting in extreme heterogeneity and long-tailed distributions. Such characteristics make prediction more difficult and contribute to the relatively low accuracy observed. Nevertheless, MoBLLM still demonstrates much stronger capabilities than LLM-Mob in capturing sparse mobility patterns, indicating that the main limitation arises from data sparsity and distribution complexity rather than the MoBLLM architecture itself. It demonstrates that an open-source LLM, i.e., LLaMA-3.1-8B-Instruct used in the case studies, can become an expert in different mobility prediction tasks by the proposed fine-tuning framework. Moreover, the MoBLLM shows a significant superiority in cost efficiency, mainly produced by generating instructions for training, which is about 60 times cheaper than a commercial LLM.
@@
4.6 Robustness under Different Prediction Situations
[…]
Overall, the MoBLLM model has robust performance in different prediction situations and significantly outperforms all the baseline models. The deep learning based models have huge performance reduction in PI-ORI and PI-DEST. Especially in PI-DEST, their performance reduces by over 10% compared with the performance in HK-DEST, while the LLM-based models only have a margin reduction of about 3%. The probable reason is that the PI datasets are much different from the HK-ORI and HK-DEST used for training the deep learning models. Deep learning models have low transferability, which leads to significant performance variation in situations where the input data is inconsistent with the training data to some extent.
We further analyze the robustness of the MoBLLM model in each prediction situation. In datasets MNC and PI that have fewer stations than the training data, the MoBLLM model performance has an average decrease by about 4% on ACC for origin prediction compared to the performance in HK-ORI. The MoBLLM model performs worse in PI than MNC for both origin and destination prediction because there exist policy interventions of price discounts that probably cause a part of passengers to change their travel behaviors. In the datasets SE and MI that have special events and incidents influencing the demands, the MoBLLM model has a margin performance reduction of about 3% for the origin prediction. It is because a small number of passengers may change their travel choices. For destination prediction on these two datasets, the MoBLLM even has better performance than the one on HK-DEST, which implies that the special events may not significantly influence the destination of existing trips. Therefore, the MoBLLM model has a good robustness to adapt to different prediction situations covering metro network change, policy disturbances, special events, and incidents.
Table 6: The descriptions of adaptation datasets.
@@ 4.6 Robustness under Different Prediction Situations
[…]
@@ 5 Conclusions
In the cases studies, six real-world mobility datasets from multiple cities are used to validate the performance of MoBLLM model. We systematically evaluate LoRA and its several advanced variants in fine-tuning to obtain the optimal MoBLLM model. Compared with the state-of-art baseline models on accuracy, the MoBLLM model averagely outperforms the best DL model (MHSA) by 50.44% and the LLM-Mob model based on commercial LLMs by 37.30%. Besides, the MoBLLM model also shows significant superiority in transferability and finance compared to the method based on commercial LLMs. In the robustness evaluation on 4 trip datasets representing different conditional scenarios, the MoBLLM shows the robust performance and outperforms the best benchmark (MHSA) by 14.67% on average. These results of case studies comprehensively demonstrate the capacity of the MoBLLM model as the foundational individual mobility prediction model.
This work has explored the application of Zero-Shot CoT prompting on the fine-tuned LLMs for individual mobility prediction. It is important to note, however, that the sequential nature of CoT reasoning introduces a vulnerability to error propagation. An incorrect fact extraction or flawed logical inference in an early step can cascade, adversely affecting all subsequent steps and the final outcome. This lack of fault tolerance presents a challenge for the deployment of fully autonomous reasoning systems. Future research should focus on integrating robustness mechanisms such as self-consistency or verification prompting.
The MoBLLM model contributes to precise and robust provision of personalized information for transportation management, e.g., crowding and disruption management. However, note that the incident logs or event information are important for prediction. Future work will involve the design of context-aware prompting strategies and the exploration of fine-tuning techniques to enhance MoBLLM’s adaptability, particularly in abnormal situations. Another promising direction involves enhancing the interpretability of LLM-generated predictions to gain deeper insights into the mechanisms governing human mobility behaviors. We also plan to extend and redevelop the proposed foundational model by evaluating its transferability to a wider spectrum of mobility prediction tasks, such as trip duration estimation. Additionally, we will explore practical applications of MoBLLM in personalized mobility research, including trip pattern analysis, to demonstrate its broader utility.
@@ Appendix A Instruction Examples
Appendix B Hallucination Cases
For the base LLaMA-3.1-8B-Instruct model, we test it on the datasets used in the case studies using the same prompts as the MoBLLM model. As shown in Figures B1 andB2, the base model often produces irrelevant content, such as Python code snippets unrelated to the prediction task, which we identify as hallucinations in this context.
Figure B1: The hallucination example for the task 1.
Figure B2: The hallucination example for the task 3.
Appendix C Hyperparameter Settings of PEFT
For PEFT methods tested in Section 4.2, we present the settings of essential hyperparameter scaling factor $\alpha$ and rank $r$ as shown in Table 1.
Table 1: The hyperparameter settings of PEFT methods.
PEFT Scaling Factor Rank
LoRA 64 64
OLoRA 16 64
EVA 1 16
PiSSA 64 64
LoftQ 16 64
LoRA+ 8 32
rsLoRA 8 32
QLoRA 8 32
Appendix D Ablation Study on Instruction Training Data
In this study, we fine-tune the LLaMA-3.1-8B-Instruct model using single-style instruction data, while keeping all other experimental conditions consistent with those of MoBLLM. We referred to this variant as MoBLLM-S. The training set incorporates the same mobility data as MoBLLM but is constructed exclusively with predefined base task prompt templates, without introducing stylistic diversity.
To assess the model’s robustness to diverse user writing styles, input prompts are generated following an approach analogous to that outlined in Section 3.4. Unlike the single-style input prompts aligned with LLM-Mob used in Section 4, the prompts in this experiment are generated by DeepSeek-Chat with a temperature of 0.3 based on the base task prompt templates to simulate a variety of writing styles. Note that employing a different LLM (DeepSeek-Chat) for generating test cases prevents leakage of styles identical to those produced by GPT-4o mini (generating training instructions for MoBLLM), thereby ensuring a fair comparison.
The results of MoBLLM-S and MoBLLM are presented in Table 1, which demonstrate the advantages of multi-style instructions for fine-tuning to achieve higher model performance.
Table 1: Ablation study on instruction training data
Model Geolife FSQ-NYC HK-ORI HK-DEST
ACC F1 ACC F1 ACC F1 ACC F1
MoBLLM-S 57.79 0.5318 40.98 0.3630 83.71 0.8362 70.46 0.6991
MoBLLM 58.37 0.5381 41.14 0.3638 84.26 0.8421 71.44 0.7093
Appendix E Model Performance for In-/Frequent Visited Locations
This study further investigates model performance on visited locations with different frequencies. First, we define the visited frequency as the ratio of the target location (ground truth) appearing in the model input (mobility sequence). To better present the stratified results, we select Geolife, FSQ-NYC and FSQ-TYK datasets and split their samples into 4-level visited frequencies as shown in Table 1. The results of Table 2 show that both LLM-Mob and MoBLLM models perform well in high-frequency samples but hardly make predictions on the target locations that have never been seen in the model input. Compared with the commercial LLM for general purpose, MoBLLM performs significantly better in the low- to medium- frequency range from 0 to 0.5, implying that fine-tuning can effectively improve the LLMs’ performance in predicting samples with such visited frequencies.
Table 1: The sample proportions (%) of 4-level visited frequencies.
Frequency Geolife FSQ-NYC FSQ-TYK
0 12.44 1.61 1.07
(0, 0.2] 32.58 66.77 86.89
(0.2, 0.5] 44.72 27.22 11.52
(0.5, 1] 10.26 4.4 0.52
Table 2: The performance comparison (ACC) under different visited frequencies.
Frequency Geolife FSQ-NYC FSQ-TYK
LLM-Mob MoBLLM LLM-Mob MoBLLM LLM-Mob MoBLLM
0 0 0.93 0 0 0 0
(0, 0.2] 29.9 44.63 12.06 25.27 10.24 22.09
(0.2, 0.5] 66.77 77.38 61.27 73.68 59.03 78.16
(0.5, 1] 90.14 87.61 90.24 92.68 94.92 96.61